# Robot Navigation with Entity-Based Collision Avoidance using Deep Reinforcement Learning

### v1

@@ Abstract
Efficient navigation in dynamic environments is crucial for autonomous robots interacting with various environmental entities, including both moving agents and static obstacles. In this study, we present a novel methodology that enhances the robot’s interaction with different types of agents and obstacles based on specific safety requirements. This approach uses information about the entity types, improving collision avoidance and ensuring safer navigation. We introduce a new reward function that penalizes the robot for collisions with different entities such as adults, bicyclists, children, and static obstacles, and additionally encourages the robot’s proximity to the goal. It also penalizes the robot for being close to entities, and the safe distance also depends on the entity type. Additionally, we propose an optimized algorithm for training and testing, which significantly accelerates train, validation, and test steps and enables training in complex environments. Comprehensive experiments conducted using simulation demonstrate that our approach consistently outperforms conventional navigation and collision avoidance methods, including state-of-the-art techniques. To sum up, this work contributes to enhancing the safety and efficiency of navigation systems for autonomous robots in dynamic, crowded environments.
Keywords Robotics  $\cdot$ Reinforcement learning  $\cdot$ Collision avoidance  $\cdot$ Deep learning  $\cdot$ Robot navigation  $\cdot$ Motion planning
@@ 1 Introduction
The development of robots, designed to assist humans and perform different tasks, has revolutionized numerous sectors, including industrial automation, manufacturing, agriculture, space, medical science, household utilities, delivery, and social services. Autonomous navigation, a key attribute of these robots, enables them to move independently, perceive the environment, and take appropriate actions to complete tasks efficiently, effectively, and safely.
@@
Recent research interest in mobile robots, particularly in the context of practical applications such as delivery, search and rescue, service, and warehouse robots has emphasized the importance of robot navigation. This field solves the problem of planning the robot’s path from the start point to the destination while avoiding other agents and obstacles in different dynamic environments. This task becomes even more challenging when the robot needs to navigate through crowded places like airports, hospitals, streets, squares, and shopping centers. Traditional navigation approaches, which often treat moving agents as static obstacles or react through short-sighted, one-step look-ahead, result in unsafe and unnatural behaviors.
The necessity for a robot to navigate in a socially compliant manner and understand the dynamic nature of its environment, especially in populated areas, is a challenging task. Furthermore, it is crucial for the robot to consider the diversity of environmental entities around it, as each entity exhibits unique characteristics, patterns of movement, safety requirements, and consequences of a collision. In this work we use the term “environmental entities", or simply entities, to collectively describe both the moving agents and the static obstacles presented in the environment. To get examples of moving agents let us consider sidewalk autonomous delivery robots. These robots are primarily used for last-mile deliveries and typically operate on pavements, courtyards, pedestrian crossings, and occasionally on parts of the road. In this case, robots interact with such moving agents as adults, children, bicyclists, scooters, vehicles, motorcycles, etc. Static obstacles in this case are curbs, trash cans, trees, bus stops, buildings, etc. For the sake of brevity, in what follows we will sometimes omit the word static, and by “obstacles" we will mean static obstacles.
@@
For robots to navigate in densely populated spaces in a socially compliant manner, it is crucial to understand the behavior of surrounding agents and obstacles and adhere to cooperative rules. To achieve that, we implement a new method, based on deep reinforcement learning, which allows the robot to navigate in dynamic environments and to interact differently with different types of agents and obstacles based on various safety requirements. We call this approach Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL).
Our main contributions can be summarized as follows:
  • We propose a new method of collision avoidance based on deep RL which can effectively utilize information about the types of entities, which helps the robot to avoid collisions and leads to safer interaction and navigation.
  • We propose a reward function that penalizes the robot differently depending on the type of agent or obstacle that the robot collides with. In our research, we consider such types of agents and obstacles: adults, bicyclists, children, and static obstacles. Apart from that, our new reward function penalizes the robot for being close to entities and the safe distance also depends on the entity. Moreover, our reward function encourages the robot not only to achieve the goal but also to be close to the goal.
  • We provide an improved algorithm for training and testing the model which significantly improved the speed of training and testing and allowed us to train the robot in complex environments.
  • We test the applicability and the effectiveness of the proposed approach in our simulator. Our experiments demonstrate that the new approach consistently outperforms standard methods of robot navigation with collision avoidance including the state-of-the-art methods.
The rest of the paper is organized as follows: in Section 2.1 we give an overview of the related work; in Section 2.2 we formulate the problem; in Section 3 the proposed EB-CADRL approach is presented, including the description of the reward function, model and training algorithm; Section 4 is dedicated to the experimental part, evaluation methodology, in this section we provide the results and ablation experiments; in Section 5 we discuss the limitations of our work and provide ideas for future work; the conclusions are made in Section 6.
2 Background
2.1 Related Work
@@
Recently extensive research has led to significant advancements in robot navigation including navigation in crowded environments.
@@
Social Force Model (SFM) Helbing and Molnár (1995) uses attractive and repulsive forces to model crowd interactions. Other approaches like Reciprocal Velocity Obstacle (RVO) and Optimal Reciprocal Collision Avoidance (ORCA) van den Berg et al. (2011) consider surrounding agents as velocity obstacles to calculate optimal collision-free velocities under reciprocal assumptions. These methods, while effective in specific scenarios, rely heavily on hand-crafted rules and often fail in real social scenarios where their assumptions are not fully met, compromising the safety of human-robot interaction and leading to the “freezing robot" problem in dense environments.
@@
Recently, learning-based methods have gained significant attention. Several studies use imitation learning to derive navigation strategies from demonstrations of desired behaviors. In studies Tai et al. (2018), Long et al. (2016) navigation policies using depth images, lidar measurements, and local maps as inputs are developed by imitating expert demonstrations. In addition to behavioral cloning, inverse reinforcement learning has been applied in Kretzschmar et al. (2016) to extract the underlying cooperation features from human data using the maximum entropy method. However, the effectiveness of these methods heavily relies on the scale and quality of the demonstrations, which can be resource-intensive and limit the policy’s quality due to human constraints.
Another approach treats robot social navigation as a Markov decision process and explores it through deep reinforcement learning, considering human safety and comfort Chen et al. (2017a). Chen et al. Chen et al. (2017b) introduce a decentralized multi-agent collision avoidance algorithm that employs a value network encoding the estimated time to the goal based on the joint state of the robot and surrounding agents. Following this, Everett et al. Everett et al. (2018) propose an enhanced algorithm using Long Short-Term Memory (LSTM) to sequentially process the states of neighboring agents, enabling navigation among an arbitrary number of agents. However, the network models in these methods are relatively simplistic and struggle to account for the crowd’s potential impact, hindering safe human-robot interaction. In contrast, Chen et al. Chen et al. (2019) present a value network that simultaneously models human-robot and human-human interactions to enhance the robot’s decision-making. Nonetheless, this method captures the spatial relationships of pedestrians but overlooks the temporal relationships of their movements, limiting the robot’s ability to predict pedestrian trajectories. In Xue et al. (2024) authors propose a deep reinforcement learning framework with a value network for mobile robot navigation that leverages spatial-temporal reasoning to understand crowd interactions, designs hazardous areas based on pedestrian speeds, and formulates a reward function to ensure pedestrian safety and comfort.
@@
2.2 Problem Formulation
We consider the task of a delivery robot navigating towards a goal in a dynamic environment populated with adults, children, bicycles, and static obstacles. This task can be formulated as a partially observable Markov decision process (POMDP). The robot must navigate in a socially compliant manner, avoiding collisions while reaching its goal. We define the state of each agent (robot or environmental entity) in terms of observable and unobservable components:
@@ 2.2 Problem Formulation
  • Unobservable states: goal position $\mathbf{g}=[g_{x},g_{y}]$, preferred speed $v_{\text{pref}}$, and heading angle $\theta$.
While acting in the environment robot knows his observable and unobservable states and also robot knows the observable states of all entities. We use the robot-centric parameterization described in Chen et al. (2017b) and Everett et al. (2018), where the robot is positioned at the origin and the x-axis is aligned with the robot’s goal direction. Denote $e_{i}$ as a type of entity $i$. The robot’s state and the $i$-th entity’s observable state at time $t$ could be represented as:
$\mathbf{s}_{t}^{r}$ $=[d_{g},v_{x},v_{y},r,v_{\text{pref}},\theta],$ (1)
@@
The optimal value function is given by Sutton and Barto (2018):
$V^{*}(\mathbf{s}_{t}^{j})=\sum_{\tilde{t}=t}^{T}\gamma^{\tilde{t}\cdot v_{\text{pref}}}R_{\tilde{t}}(\mathbf{s}_{\tilde{t}}^{j},\pi^{*}(\mathbf{s}_{\tilde{t}}^{j})),$ (4)
@@
$\pi^{*}\left(\mathbf{s}_{t}^{j}\right)=\arg\max_{\mathbf{a}_{t}\in\mathbf{A}}\biggl{[}R\left(\mathbf{s}_{t}^{j},\mathbf{a}_{t}\right)+\gamma^{\Delta t\cdot v_{\text{pref }}}\cdot$ (5)
@@
3 New Approach
In this section, we formulate our new approach which we call Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL). We propose a novel reward function. Then, we propose a model based on a neural network that utilizes the information about the types of entities. Finally, we formulate an improved version of the deep V-learning algorithm that utilizes multiprocessing and significantly improves the performance during training, validation, and testing.
@@ 3.1 Reward Function
For concise usage in the subsequent formulas we denote the entity types:
  • Adult: $A$
  • Bicycle: $B$
  • Child: $C$
  • Obstacle: $O$
@@
Let $d(e)$ be the minimum distances between the robot and each of entity types where $e\in\{A,B,C,O\}$. Let $d_{min}=\min\{d(A),d(B),d(C),d(O)\}$ be the minimum distance between the robot and all entities around the robot over the period $[t-\Delta t,t]$. If a collision with the robot occurs for entity $e$, then $d(e)$ and $d_{min}$ are less than or equal to zero.
The function $R_{t}(t)$ represents the time reward:
$R_{t}(t)=\begin{cases}1&\text{if }t<t_{good}\\ \frac{t_{max}-t}{t_{max}-t_{good}}&\text{if }t_{good}\leq t\leq t_{max}\\ 0&\text{otherwise}\end{cases}$ (6)
@@
where: $t_{good}$ is the time duration considered as “good" for reaching the goal, $t_{max}$ is the maximum allowable time to reach the goal.
@@
Let $d_{max}$ be the maximum distance to the goal, which is the Euclidean distance between the starting position of the robot at time $t=0$ and the position of the goal. Let $d_{g}(t)$ be the distance to goal at time $t$. Then proximity reward can be represented as:
$R_{p}(t)=1-\frac{d_{g}(t)}{d_{max}}$ (7)
Let $R_{c}(e)$ represent the penalties for collisions. The penalties are defined as:
$R_{c}(e)=\begin{cases}-0.5,&\text{if }e=O\\ -1.0,&\text{if }e=A\\ -1.5,&\text{if }e=B\\ -2.0,&\text{if }e=C\end{cases}$ (8)
The rationale behind the collision penalties is based on the predictability of the entity’s behavior and the severity of the consequences of a collision. Static obstacles are the most predictable and have the least severe consequences, resulting in the smallest penalty (-0.5). Adults exhibit predictable behavior with moderate collision consequences, resulting in a penalty of -1.0. Cyclists, due to their higher speed and potential for severe consequences such as falling from the bicycle, incur a higher penalty (-1.5). Children are the most unpredictable and vulnerable, leading to the most severe consequences in the event of a collision, thus resulting in the highest penalty (-2.0).
Let $d_{disc}(e)$ be the discomfort distance for entity $e$:
$d_{disc}(e)=\begin{cases}0.1,&\text{if }e=A\\ 0.2,&\text{if }e=B\\ 0.2,&\text{if }e=C\end{cases}$ (9)
Let $p_{disc}(e)$ be the discomfort penalty factor for entity $e$:
$p_{disc}(e)=\begin{cases}0.5,&\text{if }e=A\\ 1.0,&\text{if }e=B\\ 1.0,&\text{if }e=C\end{cases}$ (10)
Finally, we can construct the penalty for being close to the entity:
$c_{disc}(e)=(d(e)-d_{disc}(e))*p_{disc}(e)*\Delta t$ (11)
The robot is not penalized for proximity to static obstacles in our model, since the robot does not cause discomfort for static obstacles. Therefore, the discomfort distance $d_{disc}(O)$ and discomfort penalty factor $p_{disc}(O)$ are omitted.
Let $E_{t}$ represent the entity type of the entity which is closest to the robot at time $t$. Now, our new reward function can be represented as follows:
$R(t)=\begin{cases}R_{p}(t)&\text{if }t\geq t_{max}\text{ and }\mathbf{p}_{t}\neq\mathbf{g}\\ R_{c}(E_{t})+R_{p}(t)&\text{else if }d_{min}<0\\ 1+R_{t}(t)&\text{else if }\mathbf{p}_{t}=\mathbf{g}\\ c_{disc}(\text{C})&\text{else if }0\leq d(C)<d_{disc}(C)\\ c_{disc}(\text{B})&\text{else if }0\leq d(B)<d_{disc}(B)\\ c_{disc}(\text{A})&\text{else if }0\leq d(A)<d_{disc}(A)\\ 0&\text{otherwise }\end{cases}$ (12)
The assignment of collision penalties, discomfort penalties, and discomfort distance thresholds in our model is based on the potential risk and harm associated with the robot’s interactions with different entities. For instance, a collision with a static obstacle like a fence, wall or trash can might result in minimal damage to the robot, with no risk to other entities. Conversely, collisions with children could lead to serious injury of children. Collisions with bicyclists could lead to serious injury for both the rider and the robot and could cause the rider to fall from the bicycle, increasing the potential harm. Our model, therefore, seeks to minimize these risks by training the model to avoid such collisions and dangerous situations.
@@ 3.2 Model
Our model builds upon the architecture described in Chen et al. (2019), enhancing it for the specific task of controlling a delivery robot navigating in dynamic environments with various obstacles and pedestrians. To account for different types of entities (e.g., humans, children, bicycles), we incorporated an additional module. Initially, this module uses one-hot encoding to represent each agent type. For more advanced scenarios, one could use trainable embeddings for each agent type. These encoded features or embeddings are then concatenated with the primary features of the agents and passed into the model.
We embed the state of each entity $i$ and the map tensor $M_{i}$, along with the state of the robot, into a fixed-length vector $g_{i}$ using a multi-layer perceptron (MLP):
$g_{i}=\phi_{g}(s^{r},s^{io},M_{i};W_{g})$ (13)
where $\phi_{g}(\cdot)$ is an embedding function with ReLU activations, and $W_{g}$ are the embedding weights.
@@
Due to the significant variation in the number of surrounding entities across different scenes, our model must be able to handle a variable number of inputs while generating a fixed-size output. To achieve this, we employ a social attentive pooling module introduced in Chen et al. (2019), which learns the relative importance of each neighbor and the collective impact of the crowd in a data-driven manner.
The interaction embedding $g_{i}$ is transformed into an attention score $\alpha_{i}$ as follows:
$\begin{gathered}g_{m}=\frac{1}{n}\sum_{k=1}^{n}g_{k},\\ \alpha_{i}=\psi_{\alpha}(g_{i},g_{m};W_{\alpha})\end{gathered}$ (15)

### v2

@@ Abstract
Efficient navigation in dynamic environments is crucial for autonomous robots interacting with moving agents and static obstacles. We present a novel deep reinforcement learning approach that improves robot navigation and interaction with different types of agents and obstacles based on specific safety requirements. Our approach uses information about the entity types, improving collision avoidance and ensuring safer navigation. We introduce a new reward function that penalizes the robot for being close to or colliding with different entities such as adults, bicyclists, children, and static obstacles, while also encouraging the robot’s progress toward the goal. We propose an optimized algorithm that significantly accelerates the training, validation, and testing phases, enabling efficient learning in complex environments. Comprehensive experiments demonstrate that our approach consistently outperforms state-of-the-art navigation and collision avoidance methods.
@@ 1 Introduction
The development of robots, designed to assist humans, has revolutionized numerous sectors, including industrial automation, manufacturing, agriculture, space, medical science, household utilities, delivery, and social services [1]. Autonomous navigation, a key attribute of these robots, enables them to move independently, perceive the environment, and take appropriate actions to complete tasks successfully, safely and efficiently in terms of computational and energy resources.
@@
Recent research interest in mobile robots, particularly in the context of applications such as delivery, search and rescue, service, and warehouse robots has emphasized the importance of robot navigation [7, 13]. The core challenge is planning a path to a destination while avoiding dynamic agents and static obstacles [14], a task that is especially difficult in crowded areas. Traditional navigation methods often treat moving agents as static obstacles or use short-sighted reactions, leading to unsafe and unnatural robot behavior [9, 20].
Navigating in a socially compliant manner in populated, dynamic environments is a significant challenge [3]. A robot must account for the diversity of entities around it, as each has unique characteristics, movements and safety requirements. We use the term “environmental entities”, or simply entities, to describe both moving agents and static obstacles. For instance, a sidewalk autonomous delivery robot [10] interacts with various moving agents (e.g., adults, children, bicyclists) and static obstacles (e.g., curbs, trees).
@@
For robots to navigate in densely populated spaces in a socially compliant manner, it is crucial to understand the behavior of agents and obstacles and adhere to cooperative rules. To achieve that, we implement a new method, based on deep reinforcement learning (RL), which allows the robot to navigate in dynamic environments and interact differently with different types of agents and obstacles based on various safety requirements. We call this approach Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL).
Our main contributions are as follows. We propose a new method of collision avoidance based on deep RL which can effectively utilize information about the types of entities, which helps the robot to avoid collisions and leads to safer interaction and navigation. We propose a reward function that penalizes the robot differently depending on the type of agent or obstacle it collides with and encourages the robot to get close to the goal. Moreover, our new reward function penalizes the robot for being close to entities and the safe distance also depends on the entity. We introduce an optimized algorithm for training and testing the model which significantly accelerates the speed of training and testing and allows the robot to be trained effectively in complex environments. We test the applicability and effectiveness of the proposed approach in our simulator. Our experiments demonstrate that the new approach consistently outperforms standard methods of robot navigation with collision avoidance, including state-of-the-art methods.
2 Related Work
@@
Extensive research has led to significant advancements in robot navigation including navigation in crowded environments.
@@
Traditional methods include the Social Force Model (SFM) [11], which uses attractive and repulsive forces, and Optimal Reciprocal Collision Avoidance (ORCA) [2], which uses velocity obstacles to find collision-free paths. These approaches rely on hand-crafted rules and often fail when their assumptions are not met in real social situations.
@@
Recently, learning-based methods have gained significant attention. Several studies use imitation learning to derive navigation strategies from demonstrations of desired behaviors. In [19] the authors use generative adversarial imitation learning to develop a navigation policy that operates directly on raw depth images. In [12] maximum entropy inverse RL has been applied to extract underlying cooperation features from demonstrated pedestrian trajectories. However, the effectiveness of these methods heavily relies on the scale and quality of the demonstrations, which can be resource-intensive and limit the policy’s quality.
Deep reinforcement learning (DRL) is another prominent approach, treating social navigation as a Markov decision process [5]. Early DRL methods like CADRL [6] used value networks based on agents’ joint states, later enhanced by LSTM-RL [8] to handle a variable number of agents. These methods, however, exhibited limitations in modeling complex crowd interactions. SARL [4] improved on this by modeling both human-robot and human-human interactions, but it focused on spatial relationships while neglecting temporal dynamics. More recently, [21] proposed a DRL framework using spatial-temporal reasoning to better understand crowd behavior and ensure safety. However, none of the aforementioned approaches addresses the challenge of navigating in environments populated by different types of entities.
@@
3 Problem Formulation
We address the problem of autonomous robot navigation towards a goal in dynamic environments populated with various entities, including dynamic agents and static obstacles. This problem can be formulated as a Markov decision process (MDP). The robot must navigate in a socially compliant manner, avoiding collisions while reaching its goal. We define the state of each agent (robot or environmental entity) in terms of observable and unobservable components:
@@ 3 Problem Formulation
  • Unobservable states: goal position $\mathbf{g}=[g_{x},g_{y}]$, preferred speed $v_{\text{\text{pref}}}$, and heading angle $\theta$.
While acting in the environment, the robot knows its observable and unobservable states, as well as the observable states of all entities. We use the robot-centric parameterization described in [6] and [8], where the robot is positioned at the origin and the x-axis is aligned with the robot’s goal direction. Denote $e_{i}$ as a type of entity $i$. The robot’s state and the $i$-th entity’s observable state at time $t$ could be represented as:
$\mathbf{s}_{t}^{r}$ $=[d_{g},v_{x},v_{y},r,v_{\text{\text{pref}}},\theta],$ (1)
@@
The optimal value function is given by [18]:
$V^{*}(\mathbf{s}_{t}^{j})=\sum_{\tilde{t}=t}^{T}\gamma^{\tilde{t}\cdot v_{\text{\text{pref}}}}R_{\tilde{t}}(\mathbf{s}_{\tilde{t}}^{j},\pi^{*}(\mathbf{s}_{\tilde{t}}^{j})),$ (4)
@@
[…]
@@
4 Proposed Approach
In this section, we formulate our new approach which we call Entity-Based Collision Avoidance using Deep Reinforcement Learning (EB-CADRL).
@@ 4.1 Reward Function
Assume the robot operates in an environment with entities of $N$ different types. Let $\mathcal{E}=\{e_{1},e_{2},\ldots,e_{N}\}$ represent the set of entity types. Define $d(e_{i},t)$ as the minimum distance between the robot and any entity of type $e_{i}\in\mathcal{E}$ over the period $t-\Delta t,t]$, where $i=1,\ldots,N$.
@@
We define $d_{\text{min}}(t)=\min_{i\in\{1,\dots,N\}}d(e_{i},t)$ as the minimum distance between the robot and all entities around it over the period $t-\Delta t,t]$. If a collision occurs with an entity of type $e_{i}$, then $d(e_{i},t)$ and $d_{\text{min}}(t)$ are less than or equal to zero.
The function $R_{\text{time}}(t)$ represents the time reward:
$R_{\text{time}}(t)=\begin{cases}1&\text{if }t<t_{\text{pref}}\\ \frac{t_{\text{max}}-t}{t_{\text{max}}-t_{\text{pref}}}&\text{if }t_{\text{pref}}\leq t\leq t_{\text{max}}\\ 0&\text{otherwise}\end{cases}$ (6)
@@
where $t_{\text{pref}}$ is the preferred time duration for reaching the goal, $t_{\text{max}}$ is the maximum allowable time to reach the goal.
@@
Let $d_{\text{max}}$ be the maximum distance to the goal, which is the Euclidean distance between the starting position of the robot at time $t=0$ and the position of the goal. Let $d_{g}(t)$ be the distance to the goal at time $t$. Then proximity reward can be represented as:
$R_{\text{prox}}(t)=1-\frac{d_{g}(t)}{d_{\text{max}}}.$ (7)
Define $R_{\text{coll}}(e_{i})$ as the penalty associated with collisions involving an entity of type $e_{i}$.
Let $d_{\text{disc}}(e_{i})$ be the discomfort distance for entity of type $e_{i}$ and $p_{\text{disc}}(e_{i})$ be the discomfort penalty factor for entity of type $e_{i}$, $i=1,\ldots,N$. We can then construct the penalty for being close to the entity:
$c_{\text{disc}}(e_{i},t)=(d(e_{i},t)-d_{\text{disc}}(e_{i}))*p_{\text{disc}}(e_{i})*\Delta t.$ (8)
Let $i^{*}(t)\;=\;\arg\min_{i\in\{1,\dots,N\}}d(e_{i},t)$. For brevity, we use $i^{*}$ for $i^{*}(t)$. Then $e_{i^{*}}\in\mathcal{E}$ represents the entity type of the entity that was closest to the robot over the period $t-\Delta t,t]$, meaning $d_{\text{min}}(t)=d(e_{i^{*}},t)$. Now, our new reward function can be represented as:
[…]
The assignment of collision penalties, discomfort penalties, and discomfort distance thresholds in our model is based on the potential risk and harm associated with the robot’s interactions with different entities.
@@ 4.2 Model
Our model builds upon the architecture described in [4], enhancing it for the specific task of controlling a robot navigating in dynamic environments with various obstacles and pedestrians. To account for different types of entities (e.g., adults, children, bicycles), we incorporated an additional module. Initially, this module uses one-hot encoding to represent each agent type. For more advanced scenarios, one could use trainable embeddings for each agent type. These encoded features or embeddings are then concatenated with the primary features of the agents and passed into the model.
We embed the state of each entity $i$, along with the state of the robot, into a fixed-length vector $g_{i}$ using a multi-layer perceptron (MLP):
$g_{i}=\phi_{g}(s^{r},s^{io};W_{g}),$ (10)
where $\phi_{g}(\cdot)$ is an embedding function with rectified linear unit (ReLU) activations, and $W_{g}$ are the embedding weights.
@@
To handle varying numbers of entities, we use a social attentive pooling module [4] which learns the relative importance of each neighbor and the collective impact of the crowd in a data-driven manner. The interaction embedding $g_{i}$ is transformed into an attention score $\alpha_{i}$ as follows:
$g_{m}=\frac{1}{n}\sum_{k=1}^{n}g_{k},\qquad\alpha_{i}=\psi_{\alpha}(g_{i},g_{m};W_{\alpha}),$ (12)